OpenStudy (anonymous):

If "a" is a factor of 34, and "b" is a factor of 42, then the product of "a" and "b" could NOT be which of the following? A. 8 B. 17 C. 84 D. 102 E. 238

OpenStudy (cwrw238):

34 = 2 * 17 42 = 2 * 3 * 7

OpenStudy (cwrw238):

27?

OpenStudy (cwrw238):

it can't be 8 either

OpenStudy (cwrw238):

102 - this can be a product : 102 = 6 * 17

OpenStudy (cwrw238):

if we include 1 , 34 and 42 are factors - they ar distinct factors then 17 = 1 * 17 so thats an allowed product 2*42 = 84 - thats ok

OpenStudy (cwrw238):

34 * 7 = 238

OpenStudy (cwrw238):

so the answer is A (8)